# When OpenClaw Meets Hospital: Toward an Agentic Operating System for Dynamic Clinical Workflows

### 12 Mar 2026 v1

@@ 6.2 OS-Native Enforcement
6.3 Audit Trail as a Filesystem Property
@@
6.4 Comparison with Complex Agent Architectures
@@ 7 Discussion
This work makes four conceptual contributions to the design of AI systems for healthcare. First, we reframe the problem of deploying LLM agents in hospitals as an infrastructure design problem rather than a model capability problem. The central argument—that safety, memory, and coordination failures stem from architectural mismatches rather than from insufficient model intelligence—motivates an entirely different design strategy: constrain the environment rather than instruct the model. This perspective aligns agent deployment practice with established principles from operating systems research, particularly least-privilege execution and isolation as primary safety mechanisms. Second, we introduce a document-mutation coordination model in which all inter-agent communication is mediated through structured file writes with no direct messaging. This design reduces the inter-agent interface to two filesystem primitives, making the coordination protocol verifiable, auditable, and enforceable by OS-native mechanisms. The event subscription layer built on top of this model supports reactive coordination—including priority-flagged emergency escalation—without requiring agents to maintain persistent communication channels. Third, we propose progressive disclosure via manifest-guided navigation as an alternative to embedding-based memory retrieval. By representing the document hierarchy through human-readable manifest files and delegating navigation decisions to the agent’s language-reasoning capability, the architecture eliminates embedding infrastructure entirely while preserving the hierarchical and longitudinal structure of clinical records. Manifest-guided retrieval also produces inherently interpretable access traces: the sequence of manifests read and the children selected at each level constitute a human-readable record of what information the agent considered before reasoning. Fourth, we identify ad-hoc task composition as a clinically essential capability absent from existing hospital IT systems and show that the proposed architecture delivers it without bespoke engineering. Because fixed, pre-programmed hospital workflows cannot accommodate the long-tail variability of clinical practice, patients and physicians routinely encounter needs that installed systems cannot address. The agentic model resolves this by allowing the agent to reason over the medical skill library and compose novel task sequences on demand, augmented by the long-term memory that allows it to draw on longitudinal context accumulated across care episodes. This combination—dynamic skill composition plus longitudinal memory—makes the proposed system qualitatively more adaptive than any static workflow-based platform.
Several limitations of the current proposal warrant acknowledgment.
Manifest accuracy and maintenance cost.
@@
Manifest descriptions are themselves LLM-generated and are therefore susceptible to inaccuracy, omission, or temporal drift if the manifest update mechanism fails to fire on a mutation event. Although the maintenance policy is designed to be incremental, a high-frequency update environment—such as an intensive care unit generating continuous sensor data—may produce a volume of mutation events that stresses the LLM-call-per-update budget. Batching strategies and manifest staleness tolerances require further investigation.
Concurrent write consistency.
The current architecture does not specify a concurrency control protocol for the case in which two agents attempt to append pages to the same document simultaneously. While append-only file semantics on Linux prevent data corruption at the storage level, interleaved writes from concurrent agents could produce logically inconsistent document states. A lightweight conflict-resolution protocol, such as optimistic concurrency based on version counters already present in the mutation event format, is a natural extension but is not formalized here.
EHR integration.
@@
The architecture assumes that patient data originates within the document collection managed by the proposed system. In practice, hospitals operate established electronic health record (EHR) systems such as Epic or Cerner, and clinical data exists in HL7 FHIR-compliant formats that would need to be translated into the document hierarchy. The design of bidirectional synchronization between the proposed document store and existing EHR infrastructure is beyond the scope of this paper.
@@ 8 Conclusion
We have argued that deploying LLM agents in hospitals is primarily an infrastructure problem, not a model capability problem, and proposed the Agentic Operating System for Hospital as a concrete response. By reducing each agent’s interface to file read and write, the architecture delegates safety enforcement to the host OS, replaces vector retrieval with manifest-guided document navigation, and enables ad-hoc task composition across a curated medical skill library—together addressing the structural limitations of both general-purpose agent frameworks and fixed-function hospital IT systems. We offer this design as a foundation for clinical AI infrastructure that earns trust not by promising safe model behaviour, but by making unsafe behaviour structurally impossible.

### 21 Mar 2026 v2

@@ 3.4 Safety and governance
Audit trail as a filesystem property.
@@
Comparison with complex agent architectures.
@@
4 Evaluation
We conducted a targeted empirical evaluation of the page-indexed memory component. The goal was to test whether manifest-guided progressive disclosure recovers relevant evidence more reliably than flat vector retrieval on structured longitudinal clinical records.
4.1 Experimental setup
Data source and patient selection.
We evaluated on MIMIC-IV v2.2 [19], a de-identified electronic health record dataset from Beth Israel Deaconess Medical Center, supplemented with clinical notes from MIMIC-IV-Note [20] (available through PhysioNet [21]). Candidate patients were required to have at least three hospital admissions and to pass a quality floor: discharge-note coverage of admissions, diagnosis diversity unique ICD codes, at least one radiology report, and de-identification density tokens per 1 000 characters. From eligible candidates we drew a stratified random sample of 100 patients across gender, age bracket (50, 50–70, 70), and admission-count bracket (3–5, 6–9, 10).
Corpus construction.
[…]
To enable hierarchical navigation, manifests were generated bottom-up by the LLM as Markdown tables with four columns (Child, Type, Date, Scope). Admission-level manifests were produced from the admission’s metadata and the first 1 500 characters of each leaf document, with the Scope column summarizing each document’s clinical content in one sentence. Patient-level manifests were then produced from each patient’s demographics and a structured summary per admission (dates, document count, top three ICD diagnoses), with Scope summarizing the admission’s key events. The 914 manifests serve as the navigational index for the retrieval system.
Queries and gold labels.
Queries were generated from the selected cohort using an LLM-guided pipeline over the manifest hierarchy. For each patient, feasible tiers were inferred from corpus structure: Tier 1 required at least one leaf document, Tier 2 required at least one admission with two or more distinct document-type categories, and Tier 3 required at least two admissions containing documents. Each patient was then assigned one feasible tier by greedy deficit balancing. The LLM received the patient-level manifest, all admission-level manifests, and snippets from up to five leaf documents per admission (the first 3 000 characters of each, prioritized by document type and explicitly including laboratory results when available), and produced one candidate query together with provisional gold-document paths. The prompt enforced tier-specific constraints: Tier 1 queries had to be answerable from exactly one leaf document, Tier 2 queries from two to three documents within a single admission, and Tier 3 queries from documents spanning at least two admissions. Proposed paths were normalized against the on-disk corpus and discarded if unresolvable; when a tier remained below quota, additional queries were generated from patients feasible for that tier until 100 queries per tier were reached. This process yielded 300 evaluation queries in total, balanced evenly across the three difficulty tiers.
Gold-standard evidence sets and reference answers were constructed through a three-stage LLM-assisted pipeline. First, a sufficiency check determined whether the provisional gold documents were enough to answer the query. If not, an expansion step used the target patient’s admission-level manifests to identify additional documents from their scope descriptions, after which the expanded set was re-verified. Queries whose evidence sets remained insufficient were manually reviewed against the source corpus and corrected before benchmarking. Finally, a concise reference answer was generated using only the verified gold documents.
Models and parameters.
All LLM calls—manifest navigation, answer generation, gold-label production, and judging—used Kimi K2.5 through an OpenAI-compatible API from Alibaba Cloud. Manifest navigation and the LLM judge used max_tokens; answer generation for all three systems used max_tokens.
Systems compared.
All three systems operated in a patient-scoped setting: patient identity was resolved deterministically from the query text or the structured target_patient field, without an LLM call. Manifest-guided retrieval then navigated two levels of the hierarchy via LLM calls—first selecting admissions from the patient-level manifest (level 1, hereafter L1), then selecting documents from each admission-level manifest (level 2, L2)—to assemble a targeted set of leaf documents. Flat-RAG retrieval applied section-aware chunking (splitting on Markdown headings, falling back to paragraph boundaries, with a 512-token sliding window and 64-token overlap for oversized sections), embedded each chunk using text-embedding-v4 (1 536 dimensions, Alibaba Cloud), and performed document-level aggregation: candidate chunks were grouped by source document, ranked by a composite key (best chunk distance, supporting chunk count, mean top-3 distance), and the top- documents were selected. Metadata-filtered RAG (RAG-AF) applied the same embedding and aggregation pipeline but pre-filtered chunks by admission identifiers extracted from the query text before dense retrieval, falling back to patient-level filtering when no admission identifiers were found. This variant tests whether explicit structural metadata narrows the gap with manifest-guided retrieval. We report as the primary baseline; and appear in Table 3.
Metrics.
Retrieval precision and recall , where is the set of retrieved documents and the verified gold evidence set. We also report perfect recall (fraction of queries recovering all gold documents) and the mean number of retrieved documents. Answer quality was assessed by an LLM judge on a 1–5 rubric evaluating factual accuracy and completeness relative to the gold answer (2 100 total judgments: 300 queries 7 system configurations—manifest, three flat-RAG values, and three metadata-filtered-RAG values).
4.2 Retrieval performance
Table 1 summarizes the main comparison across three systems. Manifest-guided retrieval achieved the highest precision (0.779, 2.2 over RAG-AF and 3.4 over flat RAG) and the highest answer-quality score (4.45). On recall, manifest and RAG-AF were effectively tied overall (0.877 versus 0.876), but this aggregate masks a critical Tier-3 divergence analyzed below. Notably, RAG-AF achieved more perfect-recall queries than manifest (224/300 versus 214/300), reflecting its higher recall on Tier-1 and Tier-2; however, this came at substantially lower precision (0.352 versus 0.779) and higher document load (7.8 versus 5.8). Part of the precision gap is structural: both RAG variants always retrieve a fixed number of documents regardless of query complexity, whereas the manifest system adapts its retrieval set (1 document for Tier-1, for Tier-2, for Tier-3).
| System | Score | Precision | Recall | Perfect Recall | Docs Retrieved |
| Manifest | 4.45 | 0.779 | 0.877 | 214/300 | 5.8 |
| RAG-AF () | 4.32 | 0.352 | 0.876 | 224/300 | 7.8 |
| RAG () | 3.85 | 0.231 | 0.647 | 124/300 | 9.7 |
Table 2 breaks down performance by query tier, revealing three key patterns.
First, on Tier-1 and Tier-2 queries, metadata filtering closes the recall gap. RAG-AF reaches 0.960 and 0.967 recall on Tier-1 and Tier-2 respectively, exceeding manifest (0.870 and 0.913). This confirms that simple admission-level pre-filtering largely resolves the disambiguation problem for queries that explicitly reference an admission. However, manifest precision remains 5 higher on Tier-1 (0.860 versus 0.167) and 1.9 on Tier-2 (0.840 versus 0.441) because it retrieves only the documents needed.
Second, the clearest manifest advantage appears on Tier-3 longitudinal queries. When queries span multiple admissions without naming them, admission filtering barely helps (RAG 0.676 RAG-AF 0.701). Manifest recall (0.846) is 21% higher than RAG-AF, confirming that LLM-guided episode selection is essential for longitudinal queries.
Third, 13 of the 100 Tier-1 queries asked for the chief complaint (gold: discharge_summary.md); the navigator selected triage.md instead—a plausible but incorrect choice since both documents contain chief-complaint information. This is a navigator precision limitation rather than a structural failure and is discussed further in Section 5.
| Tier | System | Score | Precision | Recall | Perfect Recall | Docs |
| 1 | Manifest | 4.76 | 0.860 | 0.870 | 87/100 | 1.0 |
| RAG-AF () | 4.72 | 0.167 | 0.960 | 96/100 | 6.6 | |
| RAG () | 3.88 | 0.069 | 0.640 | 64/100 | 9.7 | |
| 2 | Manifest | 4.47 | 0.840 | 0.913 | 74/100 | 3.5 |
| RAG-AF () | 4.28 | 0.441 | 0.967 | 91/100 | 7.0 | |
| RAG () | 3.88 | 0.193 | 0.626 | 30/100 | 9.7 | |
| 3 | Manifest | 4.11 | 0.638 | 0.846 | 53/100 | 13.0 |
| RAG-AF () | 3.95 | 0.449 | 0.701 | 37/100 | 9.6 | |
| RAG () | 3.80 | 0.431 | 0.676 | 30/100 | 9.7 |
4.3 Sensitivity and ablation
Table 3 shows -sensitivity for both RAG variants. Even at , RAG-AF surpasses manifest recall (0.912 versus 0.877)—but at the cost of nearly double the documents (9.8 versus 5.8) and less than half the precision (0.305 versus 0.779). Manifest matches RAG-AF at on recall (0.877 versus 0.876) with 2.2 higher precision and fewer documents. Increasing further yields diminishing returns for RAG-AF while degrading precision.
| System | Precision | Recall | Perfect Recall | Docs Retrieved | Score |
| RAG () | 0.291 | 0.439 | 59/300 | 5.0 | 3.28 |
| RAG () | 0.231 | 0.647 | 124/300 | 9.7 | 3.85 |
| RAG () | 0.186 | 0.738 | 165/300 | 14.3 | 3.95 |
| RAG-AF () | 0.466 | 0.768 | 179/300 | 4.6 | 4.11 |
| RAG-AF () | 0.352 | 0.876 | 224/300 | 7.8 | 4.32 |
| RAG-AF () | 0.305 | 0.912 | 244/300 | 9.8 | 4.33 |
| Manifest | 0.779 | 0.877 | 214/300 | 5.8 | 4.45 |
To characterize the contribution of each structural layer, we constructed a four-level ablation spectrum (Table 4) ranging from pure dense retrieval to full manifest-guided navigation.
Table 4 reveals the contribution of each structural layer. Adding explicit admission metadata (RAG RAG-AF) is the single largest recall improvement (0.229), but it only works when the query text contains admission identifiers (all 100 Tier-1, all 100 Tier-2, but only 21 of 100 Tier-3 queries). LLM-based episode selection (RAG-AF Manifest L1) pushes recall to 0.977 by handling Tier-3 queries that lack explicit metadata, but precision collapses to 0.238 because all documents within selected admissions are loaded. Document-level filtering (Manifest L1 L1+L2) recovers precision to 0.779—a 3.3 improvement—with a modest recall trade-off (0.100). This confirms that the document-level manifest is the primary source of the precision advantage.
| System | Structural awareness | Precision | Recall | Perfect Recall | Docs |
| RAG () | Dense retrieval only | 0.231 | 0.647 | 124/300 | 9.7 |
| RAG-AF () | + admission metadata | 0.352 | 0.876 | 224/300 | 7.8 |
| Manifest L1 only | + LLM episode selection | 0.238 | 0.977 | 276/300 | 22.2 |
| Manifest L1+L2 | + LLM document selection | 0.779 | 0.877 | 214/300 | 5.8 |
4.4 Qualitative example
The manifest-based approach also produced directly interpretable retrieval traces. For query Q289, a Tier-3 longitudinal query asking how a patient’s lactate levels changed from an initial craniotomy admission to a later readmission for surgical site infection, the system resolved the patient from the query text, selected the two relevant admissions, and retrieved exactly two documents: lab_blood_gas.md from each admission. This trace achieved both precision and recall of 1.0. The resulting answer cited specific lactate values (2.2–3.0 mmol/L during the craniotomy admission versus 0.9 mmol/L at readmission), demonstrating that the manifest hierarchy enables clinically precise longitudinal reasoning with minimal document retrieval.
| Field | Value |
| Query | How did S19768128’s lactate levels change from the initial craniotomy admission to the later readmission for surgical site infection? |
| Selected patient | S19768128 (resolved from query text) |
| Selected admissions | hadm_20298053, hadm_20946200 |
| Selected documents | lab_blood_gas (from both admissions) |
| Retrieved files |
S19768128/hadm_20298053/lab_blood_gas.md
S19768128/hadm_20946200/lab_blood_gas.md |
| Trace quality | Precision = 1.0; Recall = 1.0 |
@@ 5 Discussion
This work makes four conceptual contributions to the design of AI systems for healthcare. We complement these architectural claims with an empirical validation of the page-indexed memory component on de-identified clinical records from MIMIC-IV. Because the benchmark evaluates document retrieval on a single institution’s records using a single LLM family, it should be interpreted as targeted evidence for the retrieval architecture rather than as end-to-end validation of a full hospital operating system. Accordingly, the evaluation supports stronger claims about retrieval quality, longitudinal evidence recovery, and trace interpretability than about OS-level safety enforcement or real-world clinical deployment.
First, we reframe the problem of deploying LLM agents in hospitals as an infrastructure design problem rather than a model-capability problem. The central argument—that failures of safety, memory, and coordination arise from architectural mismatch rather than insufficient model intelligence—motivates a different design strategy: constrain the environment rather than rely on the model alone. This perspective aligns agent deployment with established principles from operating-systems research, particularly least-privilege execution and isolation as primary safety mechanisms.
Second, we introduce a document-mutation coordination model in which all inter-agent communication is mediated through structured file writes, without direct messaging. This design reduces the inter-agent interface to two filesystem primitives, making the coordination protocol verifiable, auditable, and enforceable through OS-native mechanisms. The event-subscription layer built on top of this model supports reactive coordination, including priority-flagged emergency escalation, without requiring agents to maintain persistent communication channels.
Third, we propose progressive disclosure via manifest-guided navigation as an alternative to embedding-based memory retrieval. By representing the document hierarchy through human-readable manifest files and delegating navigation decisions to the agent’s language-reasoning capability, the architecture eliminates embedding infrastructure while preserving the hierarchical and longitudinal structure of clinical records. Manifest-guided retrieval also yields inherently interpretable access traces: the sequence of manifests read and the children selected at each level constitutes a human-readable record of the information considered before reasoning. In our MIMIC-IV evaluation, manifest-guided retrieval achieved a 2.2 precision advantage over the strongest baseline (metadata-filtered RAG) while matching its overall recall (0.877 versus 0.876) and retrieving 26% fewer documents (5.8 versus 7.8), providing concrete evidence that this navigation strategy scales to real clinical records with hundreds of documents per patient.
Fourth, we identify ad hoc task composition as a clinically essential capability that is absent from existing hospital IT systems and show that the proposed architecture can provide it without bespoke engineering for each use case. Because fixed, pre-programmed workflows cannot accommodate the long-tail variability of clinical practice, patients and physicians routinely encounter needs that installed systems cannot address. The agentic model responds by allowing the agent to reason over the medical skill library and compose novel task sequences on demand, augmented by long-term memory that draws on context accumulated across care episodes. This combination of dynamic skill composition and longitudinal memory makes the proposed system qualitatively more adaptive than static workflow-based platforms.
@@
Several limitations of the current paper warrant acknowledgment. First, manifest descriptions are themselves LLM-generated and are therefore susceptible to inaccuracy, omission, or temporal drift if the manifest update mechanism fails to fire on a mutation event. Although the maintenance policy is designed to be incremental, a high-frequency update environment—such as an intensive care unit generating continuous sensor data—may produce a volume of mutation events that stresses the LLM-call-per-update budget. Batching strategies and manifest staleness tolerances require further investigation.
Second, the concurrency-control policy is deliberately conservative. The current design serializes write-producing reasoning cycles at the patient level through short-lived filesystem leases, which is straightforward to implement and audit on Linux but may reduce throughput when many agents contend for the same high-acuity patient record. In practice, this trade-off is acceptable for the clinical setting we target because the critical safety requirement is to preserve a coherent patient state during each agent’s read–reason–write interval, and the duration of that interval is expected to be short relative to the surrounding care workflow. Nevertheless, future work should evaluate whether finer-grained locking, non-blocking append queues, or transaction-style batching can preserve the same safety guarantees while improving throughput in unusually high-frequency update environments such as intensive care monitoring.
@@
Third, the architecture assumes that patient data originates within the document collection managed by the proposed system. In practice, hospitals operate established electronic health record (EHR) systems such as Epic or Cerner, and clinical data exists in HL7 FHIR-compliant formats that would need to be translated into the document hierarchy. The design of bidirectional synchronization between the proposed document store and existing EHR infrastructure is beyond the scope of this paper.
Fourth, the evaluation uses a single LLM (Kimi K2.5) for all roles—navigation, answer generation, gold-label production, and judging—which may introduce systematic bias. The LLM judge scores reflect quality assessment by the same model family that generated the answers. Cross-model evaluation and human clinical expert review remain necessary for robust validation.
Fifth, the manifest system’s admission-level navigation granularity introduces a structural precision–recall trade-off: when a broad query activates an entire admission, sibling documents not needed for the specific query are loaded alongside relevant ones. This over-retrieval was most pronounced on tier-3 longitudinal queries (precision 0.638 versus tier-1 precision 0.860). Finer-grained navigation, such as document-type filtering within admissions, could mitigate this effect.
We tested this directly by adding admission-level metadata filtering to the RAG baseline (RAG-AF). On tier-1 and tier-2 queries where admission identifiers appear in the query text, metadata-filtered RAG closed the recall gap entirely (0.960 and 0.967 versus manifest 0.870 and 0.913) while narrowing the precision gap. However, on tier-3 longitudinal queries—where admission identifiers are typically absent—metadata filtering provided minimal improvement (0.701 versus flat RAG 0.676), and manifest recall (0.846) remained substantially higher. This confirms that the value of hierarchical navigation increases with query complexity: simple metadata filtering largely closes the gap on single-admission lookups, but cross-admission reasoning requires the kind of LLM-guided episode selection that the manifest hierarchy provides.
Our error analysis revealed two systematic patterns that illuminate the strengths and limitations of each approach. First, embedding-based retrieval systematically failed to disambiguate documents sharing vocabulary but differing by temporal or structural context. Of the 46 queries where flat RAG at achieved zero recall, 36 were tier-1 queries—typically targeting a specific document from one admission when the patient had many structurally similar documents across admissions. This pattern is consistent with our metadata-filtering experiment: RAG-AF, which pre-filters by admission identifier, reduced tier-1 zero-recall failures from 36 to 4, confirming that the disambiguation failure is fundamentally a metadata problem, not an embedding quality problem. Manifest-guided retrieval largely avoids this failure mode because it resolves the admission before selecting documents within it, though 13 tier-1 queries still failed due to document-type ambiguity within the correct admission.
Second, the manifest system’s over-retrieval on tier-3 queries followed a predictable pattern. Across 100 tier-3 queries, the system retrieved 708 non-gold documents (mean 7.08 per query), with three extreme outliers exceeding 60 extra documents each for patients with many admissions. Despite this over-retrieval, manifest tier-3 precision (0.638) still exceeded RAG-AF at (0.449).
Third, the 13 tier-1 manifest failures noted above reveal a limitation of LLM-based navigation: when two document types cover similar content, the navigator may select the wrong one. All 13 involved chief-complaint queries where the gold document was discharge_summary.md but the navigator chose triage.md—both of which contain chief-complaint information. Possible mitigation includes query-aware document ranking within the selected admission.
@@ 6 Conclusion
We argue that deploying LLM agents in hospitals is primarily an infrastructure problem rather than a model-capability problem, and we propose the Agentic Operating System for Hospitals as a concrete response. By reducing each agent’s interface to file read and file write, the architecture delegates safety enforcement to the host operating system, replaces vector retrieval with manifest-guided document navigation, and enables ad hoc task composition across a curated medical skill library. Together, these design choices address structural limitations shared by both general-purpose agent frameworks and fixed-function hospital IT systems. In an evaluation on 100 MIMIC-IV patient records with 300 stratified clinical queries (100 per tier), manifest-guided retrieval achieved 2.2 higher precision than a metadata-filtered RAG baseline (0.779 versus 0.352) while matching its overall recall, and 3.4 higher precision than flat RAG (0.779 versus 0.231). Even against the metadata-filtered baseline, manifest retrieval achieved 21% higher recall on longitudinal queries (0.846 versus 0.701), with its clearest advantages observed on multi-admission queries where simple metadata filtering is insufficient. We offer this design as a foundation for clinical AI infrastructure that earns trust not by assuming safe model behavior, but by making unsafe behavior structurally difficult.
Author Contributions
W.Y. conceived the study, developed the architecture, and wrote the manuscript. H.Q. conducted the experiments and implemented the evaluation pipeline. B.Z. and C.L. contributed clinical expertise and reviewed the clinical workflow and safety considerations. Z.H. assisted with literature review and manuscript revision. X.F., R.Y., and J.D. supervised the project and revised the manuscript. All authors reviewed and approved the final manuscript.
Competing Interests
The authors declare no competing interests.
Data Availability
The evaluation was conducted on the MIMIC-IV dataset, version 2.2 [19], with clinical notes from MIMIC-IV-Note [20], available through PhysioNet [21] (https://physionet.org/content/mimiciv/2.2/) under a credentialed data use agreement. The evaluation pipeline code is available at https://github.com/[repository-placeholder].